A monochromatic laser is exciting hydrogen atoms from the n=2 state to the n=5 state

PART A: What is the longest wavelength that is observed?

PART B: What is the shortest wavelength observed?
Answer #1

RH=1.096776×10 /m The maximum wavelength for transistion n2 = 5 to M = 4 -(1096776 x101m)| -副 = 4052 nm The minimum waveleng